Capacitance of 3uf, 6uf, and 12uf are connected in series across a 350V supply.
Calculate 1: the equivalent circuit capacitance
2: the charge on each capacitor
3: the potential difference across each capacitor
1
Expert's answer
2020-11-12T09:34:47-0500

1) So the equivalent capacity is "\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}=\frac{1}{3}+\frac{1}{6}+\frac{1}{12}=\frac{7}{12}" , hence "C \approx1. 71 \mu F"

2) Since the charges on the capacitors are equal when connected in series, "U= \frac{q}{C}"

"q=U\times C=350\times 1.71= 598.5 \mu""C"

3) "U_1=\frac{q}{C_1}=\frac{598.5 }{3}=199.5 V"

"U_2=\frac{598.5 }{6}=99.75 V"

"U_3=\frac{598.5 }{12}=49.875 V"